OpenStudy (anonymous):

if 9 + 7 = 16, and 8 + 8 = 16, then 9 + 7 = 8 + 8 which property?

OpenStudy (anonymous):

symmetric?

OpenStudy (anonymous):

Could you try to explain more what your question is? If the question is '9+7=8+8' then you are correct.

OpenStudy (anonymous):

i'm asking what the property of "if 9 + 7 = 16, and 8 + 8 = 16, then 9 + 7 = 8 + 8" <== the geometry phrase is.

OpenStudy (anonymous):

Aha, that makes it more clear, but I apologize I don't know the answer.

OpenStudy (anonymous):

Try asking in /Mathematics

OpenStudy (anonymous):

oh. ahah woops:D